OpenStudy (anonymous):

If the circumference of the circle below is 66 cm, what is the length of XY , in centimeters? Do not include units in your answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

Chord lenth, x = 2 x r x sin( a/2r ) where r = radius a is the arc length of XY

OpenStudy (anonymous):

wouldnt it just be 60/360 = 1/6 and then 66 * (1/6) = 11